Sub 26
Ques. 13

Express the logarithm in terms of log2M and log2N:

log2(MN)4
1
Expert's answer
2012-05-15T09:38:40-0400
log2(MN)^4=log2(M)^4+log2(N)^4=4log2(M)+4log2(N)